Find the general/particular solution of the following Differential Equations


(Non-Exact D.E)


(x + 4y³)dy - ydx =0


1
Expert's answer
2021-09-10T12:07:16-0400

Let us find the general solution of the differential equation "(x + 4y^3)dy - ydx =0" which is equivalent to "- yx' +x + 4y^3=0." Note that "y=0" is a solution. Let us divide both part of the last differential equation by "-y^2." Then we get "\frac{1}{y}x' -\frac{1}{y^2}x =4y." It follows that "(\frac{1}{y}x)' = 4y," and hence "\frac{1}{y}x=2y^2+C." We conclude that the general solution is "x=2y^3+Cy" or "y=0."